Calculate the percentage for the number of students who received A's, B's, and C's
Total number of students = 50
13 earned A's
33 earned B's
4 earned C's

Answer by jim_thompson5910(35256)

I'll do the A students.


Percentage of A students = 100*(Number of A students/Number of students total) = 100*(13/50) = 100*0.26 = 26%


So 26% of the 50 students earned A's